The average age of father and his son is 54 years and after 10 years the ratio of their ages will be
3 : 1. So find the present age of father.
(1) 82 years (2) 84 years (3) 86 years (4) 89 years

Answer by macston(5194)

.
F=father's age; S=son's age
.
(F+S)/2=54yrs
F+S=108yrs
S=108yrs-F
.
F+10yrs=3(S+10yrs)
F+10yrs=3(108yrs-F+10yrs)
F+10yrs=3(118yrs-F)
F+10yrs=354yrs-3F
4F=344yrs
F=86yrs
ANSWER: Father's present age is (c) 86 years.
.
CHECK:
.
S=108yrs-86yrs=22yrs
Son's age is 22 years.
.
F+10yrs=3(S+10yrs)
86yrs+10yrs=3(22yrs+10yrs)
96yrs=3(32yrs)
96yrs=96yrs